# When is the Carathéodory metric a Kähler metric?

### v1

@@ 2 Proof of the Main Result
is algebraic ($z$ is fixed), we obtain that the holomorphic sectional curvature $H(z;v)$ is equal to $-2(L(\Omega))^{2}$ for all non-zero vector $v\in T_{z}(\Omega)$. This is true for all $z\in\Omega$ such that $\delta_{\Omega}(z)<\epsilon$ for a sufficiently small $\epsilon$. Then, the holomorphic sectional curvature of the Bergman metric $B_{\Omega}$ is equal to $-2(L(\Omega))^{2}$ on $\Omega\setminus\Sigma$, where $\Sigma$ is defined by (2.2). Since the Bergman metric is complete on $\Omega$ whose boundary is $C^{2}$, by Theorem 2.3, the Bergman metric has constant negative holomorphic sectional curvature on $\Omega$. Then, Lu’s theorem in [31] implies that $\Omega$ is biholomorphic to a ball.
@@
The research of the first author was supported by an AMS-Simons travel grant.

### v2

@@ 3 Local-to-global principle for the Bergman metric
is algebraic ($z$ is fixed), we obtain that the holomorphic sectional curvature $H(z;v)$ is equal to $-2(L(\Omega))^{2}$ for all non-zero vector $v\in T_{z}(\Omega)$. This is true for all $z\in\Omega$ such that $\delta_{\Omega}(z)<\epsilon$ for a sufficiently small $\epsilon$. Then, the holomorphic sectional curvature of the Bergman metric $B^{2}_{\Omega}$ is equal to $-2(L(\Omega))^{2}$ on $\Omega\setminus\Sigma$, where $\Sigma$ is defined by (2.1). Since the Bergman metric $(\Omega,B^{2}_{\Omega})$ is complete, Theorem 1.4 yields that $\Omega$ must be biholomorphic to $\mathbb{B}^{n}$.
@@
The research of the first author was partially supported by an AMS-Simons travel grant and the NSF grant DMS-2103608.
Acknowledgements
The authors sincerely thank Łukasz Kosiński, Bernhard Lamel, Damin Wu and Andrew Zimmer for their comments. The first author thanks Song-Ying Li, Ming Xiao, and Hang Xu for stimulating discussions.